# Effective synchronization amid noise-induced chaos

### v2

@@ IV.1 Proof of effective synchronization
We now test the extent to which agent $\mathrm{A}$}’s $\varphi_{{\mathrm{f}},{\mathrm{A}}}^{(k)}$ and $\mathrm{B}$}’s $\varphi_{{\mathrm{f}},{\mathrm{B}}}^{(k)}$ agree. We use an extensive simulation of the phase map of Eq. (2) with $\Lambda=0.141$ and $N=500$. We extend the simulation depicted in Fig. 2 to a large number of kicks, $k=10^{4}$. For every $k$, using the procedure of Appendix A.2 (Eq. (7)), we find $\varphi^{(k)}_{{\mathrm{f}},{\mathrm{A}}}$ and $\varphi^{(k)}_{{\mathrm{f}},{\mathrm{B}}}$. To quantify the success of our scheme for identifying a common fiducial phase, we consider the statistics of the deviation $\Delta\varphi_{\mathrm{f}}=\varphi_{{\mathrm{f}},{\mathrm{A}}}-\varphi_{{\mathrm{f}},{\mathrm{B}}}$. There are indeed $4,987$ positive and $5,013$ negative $\Delta\varphi_{\mathrm{f}}$ values, both exhibiting identical histograms (the KLD among the two $\ln(\pm\Delta\varphi_{\mathrm{f}})$ distributions is $\mathcal{O}(10^{-3})$), so we only show $|\Delta\varphi_{\mathrm{f}}|$ henceforth. Likewise, in light of the convergence we observe in Sec. III, we shall only plot $S=(S_{\mathrm{A}}+S_{\mathrm{B}})/2$, where $S_{\mathrm{A}}$ and $S_{\mathrm{B}}$ only differ by numerical entropy estimation error $\sim 0.2$ for large $k\gg K_{\mathrm{m}}$.
Though $\Delta\varphi_{\mathrm{f}}$ varies widely between kicks, this variability stems from the erratically varying entropies of the distributions $q^{(k)}$ (Fig. 3(b)). In Fig. 5(a), we plot the scaled deviation $\ell\equiv\sqrt{N}|\Delta\varphi_{\mathrm{f}}|/e^{S}$ versus $S$, where the color intensity depicts the joint probability density for $(S,\log_{10}\ell)$. We see that for all $S$’s, most $|\Delta\varphi_{\mathrm{f}}|$’s lie close to $e^{S}/\sqrt{N}$ (see the $\ell=1$ line). The deviations from the peak appear to fall off roughly as a Gaussian — we further resolve this in Fig. 5(b). There, we plot the marginal cumulative distribution of the rescaled deviation $\ell$ irrespective of $S$, highlighting the abundance of small deviations as seen qualitatively in Fig. 5(a). The inset shows the marginal cumulative distribution of the entropy $S$ irrespective of $\ell$. The raw data for Fig. 5 — the $10^{4}$ instances of deviations $|\Delta\varphi_{\mathrm{f}}|$ versus $S$ — is shown in SI Appendix, Fig. S8(a).
@@
To conclude, this section has provided concrete evidence supporting the proposal made in the Introduction. Identical nonlinear oscillators exposed to identical noise can enable independent agents to act in concert and perform time-based communications as though their clocks were synchronized, even though they are not and their phases are only statistically determined.
@@ V Discussion
In this paper, we showed that noise-induced synchronization can be extended beyond its recognized limits. Our study demonstrated effective synchronization under noise-driven dynamics with positive Lyapunov exponents, which precludes conventional noise-induced synchronization. This effective synchronization arose from two key features of our dynamics. First, the statistical distribution of phases at a given moment becomes independent of the initial distribution for arbitrary realization of waiting times, as noted in Sec. III. Thus, the distributions seen by two independent agents become equal to each other while varying (together) strongly with more kicks. Second, concentrating on dynamics whose Lyapunov exponent is small in magnitude, we found in Sec. IV that the prevalence of low-entropy distributions $q(\varphi)$ permits accurate effective synchronization.
In this section, we assess the potential impact of these results. First, we note the narrow though significant scope of our explicit study. We argue that this generalized synchronization should occur generically for oscillators subjected to non-synchronizing noise, despite possible degrading effects we ignored. We then discuss the power-law trade-off we observed between the desirable rapid convergence and the obtainable precision of the generalized synchronization. Finally, we consider how generalized synchronization might be relevant for explaining forms of cooperative behavior in living systems.
@@ V.1 Limitations
Our demonstration of effective synchronization was made in a narrow context. Among the various types of noise treated in the literature [18, 25], we considered only impulsive noise whose effect could be described by a phase map [22, 24]. Furthermore, the bulk of our study used a simple class of such maps — a cubic polynomials; other phase maps may not culminate in low-entropy distributions or have a meaningful fiducial phase. These limitations mean that our example is not immediately applicable to realistic conditions. Still, we may argue that the effect is somewhat general and robust. The effective synchronization persists under continuous variation of the phase map $\psi(\varphi)$ through cubic and quintic phase maps and over a range of Lyapunov exponents $\Lambda$. Further, we recover the effective phase $\varphi_{\mathrm{f}}$ using standard-precision computations and conventional sampling methods. This is not a delicate effect analogous to time-reversing a chaotic trajectory. Moreover, all of our results in Sec. IV show a common variation with the distribution entropy $S$, $\Delta\varphi_{\mathrm{f}}\sim e^{S}/\sqrt{N}$, under a range of underlying dynamics that gave rise to the distribution. Under general noise, one may define the Lyapunov exponent [18, 21], the distribution of $\varphi$ values $q(\varphi)$ at a given time, the notion of independence of this $q(\varphi)$ on initial conditions, and the determination of a fiducial phase $\varphi_{\mathrm{f}}$ from a given $q(\varphi)$. Thus, the synchronization we found in our narrow context should plausibly have a counterpart for clocks subjected to more general forms of noise, such as sensory neural networks [31].
@@
In addition, we did not address counter-effects that oppose synchronization, such as direct interactions [47, 48, 49], the presence of an additional degrading noise that acts independently on each oscillator [20, 19], and inevitable differences between the two agents’ oscillators [19]. These must be included to reliably model, e.g., ecological systems where spatial variations in the environment or inherent differences in the individuals are of interest. We comment, however, that these counter-effects plague standard noise-induced synchronization ($\Lambda<0$) just as they would our effective synchronization ($\Lambda>0$). For $\Lambda<0$, it was shown that these effects ultimately lead to an asymptotic $q^{(k)}(\varphi)$ whose nonzero width is determined by the magnitude of the corresponding counter-effect [19, 20] (as opposed to an ever-decreasing width of $q^{(k)}(\varphi)$ for standard noise-induced synchronization). We expect a similar conclusion for $\Lambda>0$, whereby the scaling $\sim e^{S}/N^{1/2}$ might worsen as a function of degrading noise’s magnitude. This will be explored in a future work.
@@
Although our prescription for identifying $\varphi_{\mathrm{f}}$ is somewhat ungainly, the result is about as precise as one might hope. The statistical fluctuations of $\Delta\varphi_{\mathrm{f}}$ appear consistent with an $N$-sample average of a Gaussian distribution whose variance is of order $N\langle\Delta\varphi_{\mathrm{f}}^{2}\rangle$, where $\langle\cdots\rangle$ denotes the average over samples. The entropy of the $\Delta\varphi_{\mathrm{f}}$ distribution is thus of order $\ln(\langle\Delta\varphi_{\mathrm{f}}^{2}\rangle^{1/2})+\log N$ [46]. Now, the scaling shown in Fig. 5 amounts to saying $\ln(\langle\Delta\varphi_{\mathrm{f}}^{2}\rangle^{1/2})\simeq S-\log N$, i.e., $S$ is comparable to the entropy of $\Delta\varphi_{\mathrm{f}}$. Thus, a much more precise $\varphi_{\mathrm{f}}$ than the observed one would need to contain more information than the source distribution $q(\varphi)$ from which it was derived — a contradiction. Thus, we cannot expect other methods to yield qualitative improvements in precise synchronization compared to $\Delta\varphi_{\mathrm{f}}\sim e^{S}/\sqrt{N}$. [^10] One straightforward example of a fiducial phase includes the first circular mean, whose estimation error follows the central limit theorem, thus scaling as $\sigma/\sqrt{N}$ with $\sigma\sim e^{S}$ (for $\sigma\ll 1$, which are the useful distributions; see, e.g., Ref. [50] on directional statistics). We prefered to use the highest-peak position as the fiducial phase since it is more illustrative of the effect we report compared to the circular mean — that peak’s dominance over all others lies at the core of the effective synchronization. Nevertheless, one can hope for more insightful and efficient calculation methods, which might or might not involve a fiducial phase.
@@ V.2 Convergence rate scaling
The central feature of our dynamics that makes effective synchronization possible is the property of convergence; the final distribution $q(\varphi)$ depends only on the recent history of waiting times. Interestingly, we found that the amount of history required had a simple dependence on $\Lambda$; the effective number of relevant kicks, $K_{\mathrm{m}}$, followed a simple power law. This observed power-law dependence merits further discussion.
@@
In general, when an iterated map becomes independent of its initial state, the characteristic kick number for convergence $K_{\mathrm{m}}$ has little relation to $\Lambda$. $\Lambda$ characterizes the separation rate of adjacent points while $K_{\mathrm{m}}$ depends on the merger of separated regions. Nonetheless, we empirically find a systematic variation of $K_{\mathrm{m}}$ with $\Lambda$ and, even more so, that $K_{\mathrm{m}}$ diverges as $\Lambda\to 0$. Characterizing this divergence further showed an apparent near-$3/2$ power-law dependence ranging across two orders of magnitude in $K_{\mathrm{m}}$. This suggests that there is some unexplained interdependence between $K_{\mathrm{m}}$ and $\Lambda$, as we sketched qualitatively in Sec. III. However, we have no evidence for such interdependence beyond the cubic map. Understanding this divergence merits a future study.
We note that the convergence found here is quite unusual. In the paradigmatic dynamics exhibiting mixing [37], any initial condition asymptotically approaches an invariant distribution. Here, since the phase map keeps being randomly applied, there is no invariant distribution $q^{(k\to\infty)}(\varphi)$, and instead every consecutive $q^{(k)}(\varphi)$ differs stochastically from the previous. It is the $q^{(k)}$’s with a different $q^{(0)}$ that become progressively more similar with $k$ for a given $\{\beta^{(k)}\}$. Such a convergence phenomenon was shown to occur in single-particle chaotic trajectories under common noise as well [51], and of course occurs in conventional noise-induced synchronization.
For practical purposes, with a known scaling between $K_{\mathrm{m}}$ and $\Lambda$, one can readily predict the increase in $\Lambda$ needed to effect a desired reduction in $K_{\mathrm{m}}$. Then, from $S\sim 1/\Lambda$ of Ref. [24] and $\Delta\varphi_{\mathrm{f}}\sim e^{S}/\sqrt{N}$ from Sec. IV, one can infer the expected loss of precision in $\varphi_{\mathrm{f}}$.
V.3 Biological implications
We posit that this synchronization scheme could enable biological agents — cells, organelles, or gene expression networks — to cooperate in a wider range of noisy environments than previously envisioned. For example, our mechanism may be relevant to a long-standing experimental puzzle about information transmission between neurons via sequences of pulses or spikes [49]. The timing of these spikes is believed to carry that information [4, 8, 9]; thus, there needs to be some means of calibrating this timing. The present work suggests that the ambient electrical noise in the brain or variations of the sensed environment [5, 6, 31] could aid this calibration. In particular, these effects may serve as sources for common noise, which can produce a generalized notion of synchronization.

### v3

@@ IV.1 Proof of effective synchronization
Figure 5: (a) A density plot of the rescaled discrepancy, $\ell=\sqrt{N}|\Delta\varphi_{\mathrm{f}}|/e^{S}$ versus $S=(S_{\mathrm{A}}+S_{\mathrm{B}})/2$. The shade of blue represents the indicated values of the joint probability density function (PDF) to obtain $S$ and $\log_{10}\ell$, relative to its maximum. It was computed with kernel-density estimation from the $k=10^{4}$ data points shown in Appendix I (Fig. 13(a)), using a Gaussian kernel of width $0.4$ in $S$ and $0.2$ in $\log_{10}\ell$. The maximal value is $\mathrm{PDF}(-1.94,-0.0052)=0.135$, whose position is indicated by the black point. The puzzling lobe of probability at the upper right is addressed in Appendix I. (b) The marginal cumulative distribution function (CDF) of the discrepancy $\log_{10}\ell$, which was arranged into bins of size $0.25$. The dashed line is the error function of width unity; its close agreement with the data suggests that the discrepancy is predominantly normal-distributed with the expected scaling $|\Delta\varphi_{\mathrm{f}}|\sim e^{S}/\sqrt{N}$. Inset: The marginal CDF of the entropy $S$, which was arranged into bins of size $0.5$. The dashed linear line implies, up to the low- and high-entropy outliers, that the entropy is exponentially distributed [24].
We now test the extent to which agent $\mathrm{A}$}’s $\varphi_{{\mathrm{f}},{\mathrm{A}}}^{(k)}$ and $\mathrm{B}$}’s $\varphi_{{\mathrm{f}},{\mathrm{B}}}^{(k)}$ agree. We use an extensive simulation of the phase map of Eq. (2) with $\Lambda=0.141$ and $N=500$. We extend the simulation depicted in Fig. 2 to a large number of kicks, $k=10^{4}$. For every $k$, using the procedure of Appendix G (and the definition in Eq. (7)), we find $\varphi^{(k)}_{{\mathrm{f}},{\mathrm{A}}}$ and $\varphi^{(k)}_{{\mathrm{f}},{\mathrm{B}}}$. To quantify the success of our scheme for identifying a common fiducial phase, we consider the statistics of the deviation $\Delta\varphi_{\mathrm{f}}=\varphi_{{\mathrm{f}},{\mathrm{A}}}-\varphi_{{\mathrm{f}},{\mathrm{B}}}$. There are indeed $4,987$ positive and $5,013$ negative $\Delta\varphi_{\mathrm{f}}$ values, both exhibiting identical histograms (the KLD among the two $\ln(\pm\Delta\varphi_{\mathrm{f}})$ distributions is $\mathcal{O}(10^{-3})$), so we only show $|\Delta\varphi_{\mathrm{f}}|$ henceforth. Likewise, in light of the convergence we observe in Sec. III, we shall only plot $S=(S_{\mathrm{A}}+S_{\mathrm{B}})/2$, where $S_{\mathrm{A}}$ and $S_{\mathrm{B}}$ only differ by numerical entropy estimation error $\sim 0.2$ for large $k\gg K_{\mathrm{m}}$.
Though $\Delta\varphi_{\mathrm{f}}$ varies widely between kicks, this variability stems from the erratically varying entropies of the distributions $q^{(k)}$ (Fig. 3(b)). In Fig. 5(a), we plot the scaled deviation $\ell\equiv\sqrt{N}|\Delta\varphi_{\mathrm{f}}|/e^{S}$ versus $S$, where the color intensity depicts the joint probability density for $(S,\log_{10}\ell)$. We see that for all $S$’s, most $|\Delta\varphi_{\mathrm{f}}|$’s lie close to $e^{S}/\sqrt{N}$ (see the $\ell=1$ line). The deviations from the peak appear to fall off roughly as a Gaussian — we further resolve this in Fig. 5(b). There, we plot the marginal cumulative distribution of the rescaled deviation $\ell$ irrespective of $S$, highlighting the abundance of small deviations as seen qualitatively in Fig. 5(a). The inset shows the marginal cumulative distribution of the entropy $S$ irrespective of $\ell$. The raw data for Fig. 5 — the $10^{4}$ instances of deviations $|\Delta\varphi_{\mathrm{f}}|$ versus $S$ — is shown in Appendix I (Fig. 13(a)).
@@
To conclude, this section has provided concrete evidence supporting the proposal made in Sec. I. Identical nonlinear oscillators exposed to identical noise can enable independent agents to act in concert and perform time-based communications as though their clocks were synchronized, even though they are not and their phases are only statistically determined.
@@ V Discussion
In this paper, we showed that noise-induced synchronization can be extended beyond its recognized limits. Our study demonstrated effective synchronization under noise-driven dynamics with positive Lyapunov exponents, which precludes conventional noise-induced synchronization. This effective synchronization arose from two key features of our dynamics. First, the statistical distribution of phases at a given moment becomes independent of the initial distribution for any $\beta^{(k)}$ realization after a well-defined finite typical mixing-kick number $K_{\mathrm{m}}$, as noted in Sec. III. Thus, the distributions seen by two independent agents become equal to each other while varying (together) strongly with more kicks. Second, concentrating on dynamics whose Lyapunov exponent is small in magnitude, we found in Sec. IV that the prevalence of low-entropy distributions $q(\varphi)$ permits accurate effective synchronization.
With these results in hand, we outline one possible communication protocol between two remote agents. Before departing, the two agents must first agree on acceptable KLD and entropy thresholds. Then, each agent’s $N$ oscillators are used to infer those quantities. For instance, by having each agent split their oscillator population into two, they may independently estimate the KLD and decide if their initial oscillator configurations have been forgotten based on whether the KLD lies below the agreed-upon KLD threshold. Then, between successive kicks, if the independently estimated entropy lies below the agreed-upon entropy threshold, communication should take place using their independently inferred fiducial phases (both rotating on the phase circle at the same rate), where the uncertainty is known by each agent to be $\sim e^{S}/\sqrt{N}$. If an unfavorable kick has occurred (resulting in a high entropy), each agent would independently be able to decide to cease communication until the next beneficial kick. Given the phenomena we have established in our paper, the independent decisions are guaranteed to be consistent between the two agents. Thus we establish effective synchronization, wherein the individual phase samples are not synchronized, but a collective phase variable that is constructed independently out from them — the fiducial phase — is. The extent to which the fiducial phases agree varies stochastically over time, but can independently be predicted by estimating the current entropy.
Below, we assess the potential impact of these results. First, we note the narrow though significant scope of our explicit study. We argue that this generalized synchronization should occur generically for oscillators subjected to non-synchronizing noise, despite possible degrading effects we ignored. We then discuss the power-law trade-off we observed between the desirable rapid convergence and the obtainable precision of the generalized synchronization. Finally, we consider how generalized synchronization might be relevant in cryptography, quantum-scale communication, and for understanding forms of cooperative behavior in living systems.
@@ V.1 Limitations
Our demonstration of effective synchronization was made in a narrow context. Among the various types of noise treated in the literature [18, 25], we considered only impulsive noise whose effect could be described by a phase map [22, 24]. Furthermore, the bulk of our study used a simple class of such maps — a cubic polynomials; other phase maps may not culminate in low-entropy distributions or have a meaningful fiducial phase. These limitations mean that our example is not immediately applicable to realistic conditions. Still, we may argue that the effect is somewhat general and robust. The effective synchronization persists under continuous variation of the phase map $\psi(\varphi)$ through cubic and quintic phase maps and over a range of Lyapunov exponents $\Lambda$. Further, we recover the effective phase $\varphi_{\mathrm{f}}$ using standard-precision computations and conventional sampling methods. This is not a delicate effect analogous to time-reversing a chaotic trajectory. Moreover, all of our results in Sec. IV show a common variation with the distribution entropy $S$, $\Delta\varphi_{\mathrm{f}}\sim e^{S}/\sqrt{N}$, under a range of underlying dynamics that gave rise to the distribution. Under general noise, one may define the Lyapunov exponent [18, 21], the distribution of $\varphi$ values $q(\varphi)$ at a given time, the notion of independence of this $q(\varphi)$ on initial conditions, and the determination of a fiducial phase $\varphi_{\mathrm{f}}$ from a given $q(\varphi)$. Thus, the synchronization we found in our narrow context should plausibly have a counterpart for clocks subjected to more general forms of noise.
@@
In addition, we did not address counter-effects that oppose synchronization, such as direct interactions [50, 51], the presence of an additional degrading noise that acts independently on each oscillator [20, 19], and inevitable differences between the two agents’ oscillators [19]. Further, the two oscillators will in general experience random differences in arrival times of the kicks or other noise. These must all be included in order to reliably model, e.g., ecological systems, where spatial variations in the environment or inherent differences in the individuals are of interest. Nonetheless, we believe that achieving our strong-noise-induced effective synchronization in the presence of interactions or nonidentical perturbations is a realistic aim in many situations. Our belief is based in part on existing studies of conventional noise-induced synchronization ($\Lambda<0$), where the same degrading effects noted above have been studied and found manageable [19, 20]. Analogous manageable degradation should therefore be plausible for our effective synchronization ($\Lambda>0$). Explicitly, for $\Lambda<0$, it was shown that an additional, uncommon noise ultimately led to an asymptotic $q^{(k)}(\varphi)$ whose nonzero width is determined by the magnitude of the uncommon noise [19, 20]. [^10] These references do not consider the effect of timing uncertainty explicitly, though Ref. [19] did consider the effect of differing oscillator frequency, which we expect to be more damaging. We expect a similar conclusion for $\Lambda>0$, in which the scaling $\sim e^{S}/N^{1/2}$ should worsen as a function of the degrading noise’s magnitude. Furthermore, the combination of interactions and shared noise was shown to still permit synchronization in the human brain [52] and in quantum circuits [53]. These aspects will be explored in a future work.
@@
Although our prescription for identifying $\varphi_{\mathrm{f}}$ is somewhat ungainly, the result is about as precise as one might hope. The statistical fluctuations of $\Delta\varphi_{\mathrm{f}}$ appear consistent with an $N$-sample average of a Gaussian distribution whose variance is of order $N\langle\Delta\varphi_{\mathrm{f}}^{2}\rangle$, where $\langle\cdots\rangle$ denotes the average over samples. The entropy of the $\Delta\varphi_{\mathrm{f}}$ distribution is thus of order $\ln(\langle\Delta\varphi_{\mathrm{f}}^{2}\rangle^{1/2})+\log N$ [49]. Now, the scaling shown in Fig. 5 amounts to saying $\ln(\langle\Delta\varphi_{\mathrm{f}}^{2}\rangle^{1/2})\simeq S-\log N$, i.e., $S$ is comparable to the entropy of $\Delta\varphi_{\mathrm{f}}$. Thus, a much more precise $\varphi_{\mathrm{f}}$ than the observed one would need to contain more information than the source distribution $q(\varphi)$ from which it was derived — a contradiction. Thus, we cannot expect other methods to yield qualitative improvements in precise synchronization compared to $\Delta\varphi_{\mathrm{f}}\sim e^{S}/\sqrt{N}$. Nevertheless, one can hope for more insightful and efficient calculation methods, which might or might not involve a fiducial phase.
@@ V.2 Theoretical implications
The central feature of our dynamics that makes effective synchronization possible is the property of finite-time convergence; the final distribution $q(\varphi)$ depends only on the recent history of waiting times. Although rigorous proofs, based on the spreading-of-phases argument, explain why asymptotic convergence is a possible outcome [29, 30, 31, 32], the behavior of the convergence time had not been characterized before. Its predictability is central to making our scheme practically relevant in the first place. Even more so, curiously, we found that the amount of history required had a simple dependence on $\Lambda$; the effective number of relevant kicks, $K_{\mathrm{m}}$, followed a power law.
@@
In general, when an iterated map becomes independent of its initial state, the characteristic kick number for convergence $K_{\mathrm{m}}$ has little relation to $\Lambda$. $\Lambda$ characterizes the separation rate of adjacent points while $K_{\mathrm{m}}$ depends on the merger of separated regions. Nonetheless, we empirically find a systematic variation of $K_{\mathrm{m}}$ with $\Lambda$ and, even more so, that $K_{\mathrm{m}}$ diverges as $\Lambda\to 0$. Characterizing this divergence further showed an apparent near-$3/2$ power-law dependence ranging across two orders of magnitude in $K_{\mathrm{m}}$. In Sec. III, we have given one qualitative argument for this interdependence, wherein $\Lambda>0$ implies that the phase map “expands and fold” neighboring phases, which in turn is responsible for the convergence. Without the former ($\Lambda\to 0^{+}$), the latter should not occur ($K_{\mathrm{m}}\to\infty$). We emphasize, however, that we have no evidence for such interdependence beyond the cubic map; understanding these scalings merits a future study.
As a corollary, we mention the close resemblance that noise-induced synchronization bears to the concept of random attractors [40]. Just as deterministic systems described by ordinary differential equations may possess different types of attractors, random dynamical systems may admit attractors that depend on the noise realization. A specified noise realization determines the time-dependent basin of attraction and its region of support, which may differ between noise realizations and will keep evolving in time with subsequent noise. Our passage along the $\Lambda$ axis from evolution towards a (stochastic set of) synchronized single phase to a state characterized by a distribution of phases is reminiscent of known random dynamical systems at a bifurcation point. These bifurcating systems may shed light on our system, particularly our observed scaling of relaxation number of kicks $K_{\mathrm{m}}$ with the Lyapunov parameter $\Lambda$. Conversely, the effective synchronization demonstrated here may have analogs in these bifurcating systems.
V.3 Practical implications
We expect that the effective synchronization explored above may appear in a range of contexts. Here, we conjecture some of the possible utilities of common noise. For example, common noise adds to the possible ways that evolved systems such as biological organisms might exploit complex dynamics to generate subtle, adaptive behavior. Extending our strong-noise approach to account for interacting or stochastic clocks might be promising and relevant extensions of our work. They would allow testing whether noise-based encoding schemes might capture communication among, e.g., neurons such as central pattern generators [54, 55] and motion-sensitive neurons [4, 6].
Further, our mechanism might offer promising implications, for example, in cryptographic contexts [56]. We posit that two remote agents may communicate amid chaos by utilizing the ambient noise as a time-dependent encryption key, potentially making it challenging for eavesdropping parties to interfere. Decoding the messages would require noting only the recent history of noise.
Another potential application of our mechanism is to synchronization-based quantum communication systems [57, 58, 59]. With the growing recognition that noise-induced synchronization can occur also in qubit chains [26, 27], we suggest that in the regime where the noise does not entangle the edge spins, they nonetheless should be partially correlated by virtue of statistical convergence onto low-entropy distributions.